ISS Daily Summary Report – 12/20/13

Extravehicular Activity (EVA) preparation: Flight Engineer (FE)-3 Hopkins, FE-5 Mastracchio, and FE-6 Wakata performed a final tool gather and audit needed for tomorrow’s planned EVA. The crew also reviewed procedures associated with the Ammonia decontamination protocol, EVA timelines, and a briefing package for Long Duration Tie Down Tether (LDTDT) configuration.  EVA #24 hatch open is planned for  ~6:10 am on Saturday, December 21st

Robotics Onboard Trainer (ROBOT) Session 2:  Wakata completed a 2nd session on this simulator to gain proficiency in Space Station Remote Manipulator System (SSRMS) Operations associated with the ETCS Loop A Pump Module removal and replacement.

Node 1 (N1) Isolation Jumper Removal: Hopkins removed the N1 isolation jumper from N1 RS1 Secondary Power Distribution Assembly (SPDA). This activity allowed recovery of power bus Z14B if there was a DC to DC Converter Unit (DDCU) S04B or other upstream power failures.

Mobile Servicing System (MSS) Operations: Robotics Ground Controllers powered up the MSS and walked the Space Station Remote Manipulator System (SSRMS) off the Node 2 Power Data Grapple Fixture (PDGF) onto Mobile Base System (MBS) PDGF1 and configured it for the Mobile Transporter (MT) translation. They then translated the MT from Worksite 4 (WS4) to WS2. Finally, they maneuvered the SSRMS into position for the start of tomorrow’s EVA and configured the SSRMS, the Special Purpose Dexterous Manipulator (SPDM) and the MBS Payload and On-orbit Replaceable Unit (ORU) Accommodation (POA) as required to support EVAs to remove and replace the Loop A Coolant Pump Module.

External Thermal Control System (ETCS) Loop A Depress: Flight Controllers reduced the ETCS Loop A pressure to 180 psi  to help Quick Disconnect (QD) manipulation during the upcoming EVA.

Oxygen Generation Assembly (OGA) Status: Last night an OGA Rotary Separator Accumulator (RSA) fill occurred resulting in a fast shutdown and loss of USOS O2 generation and water circulation. Engineering teams believe the problem is due to a small amount of free gas in the sensor but are conducting evaluations to verify.

Waste and Hygiene Compartment (WHC) Status: The WHC pump separator failed earlier today.  The crew was able to complete a pump separator R&R and the WHC is operating nominally at this time.

Carbon Dioxide Removal Assembly (CDRA) Faults: The Node 3 CDRA experienced several faults since yesterday but has been recovered each time.  Valve ASV 101 was replaced at the end of October. There are no pristine spares on orbit.  There are two ASV valves manifested on Orbital-1 currently scheduled to launch no earlier than January 7, 2014. There are two degraded spares on orbit that were in operation but removed due to performance issues.  Teams are assessing a forward plan.

Today’s Planned Tasks

  • Morning Inspection. SM ПСС (Caution & Warning Panel) Test [Complete]
  • ISS crew and ГОГУ (RSA Flight Control Management Team) weekly conference [Complete]
  • EVA – Decontamination OBT [Complete]
  • EVA tool config [Complete]
  • Equipment Lock prep [Complete]
  • VIRTUAL. Set up, Conduct Experiment and Closeout Ops [Complete]
  • EVA systems update [Complete]
  • ROBOT startup [Complete]
  • OBSTANOVKA. Download of Obstanovka P/L Data from БСПН to RSS1 [Complete]
  • PILLE Dosimeter Prep for US EVA [Complete]
  • ROBOT OBT [Complete]
  • Spacesuit comm and telemetry configuration [Complete]
  • EMU seal lubricant application [Complete]
  • EVA tool audit [Complete]
  • BETA-08 Data Output Device (УСИ) Checkout Using Gamma-1M [Complete]
  • Orlan and БСС telemetry check. Orlan comm check and medical parameter test  [Complete]
  • СОЖ Maintenance [Complete]
  • Space Headaches questionnaire[Complete]
  • ER8 locker removal [Complete]
  • Journals [Complete]
  • EVA procedures review [Complete]
  • ROBOT procedures review [Complete]
  • RWS setup [Complete]
  • Ku Band power supply setup [Complete]
  • P/TV SSC prep P/TV VTR bypass config [Complete]
  • EVA S/G tagup [Complete]
  • Comm reconfig into initial after Orlan Comm. Check [Complete]
  • VIRTUAL. Set up, Conduct Experiment and Closeout Ops [Complete]
  • В3 Fan Screen Cleaning in DC1 [Complete]
  • OBSTANOVKA. Download of Obstanovka P/L Data from БСПН to RSS1 [Complete]
  • Consolidate Tool Caddy, Photography, Photo Downlink r/g 4351 [Complete]
  • Installation of Add-on Equipment on ORLAN-MK for EVA-37 [Complete]
  • Cleaning ВД1 and ВД2 Air Ducts in DC-1 [Complete]
  • SSRMS DOUG setup [Complete]
  • ПФ1, ПФ2 Dust Filter Replacement and В1, В2 Fan Grille Cleaning in DC1 [Complete] 

Task List

  • ARED Cylinder Evacuation
  • IMS Update
  • RS transfer
  • Roscosmos Report
  • URAGAN. Observations and Photography
  • ECON-M. Observations and Photography
  • Extravehicular Activity (EVA) Procedure Review
  • National Geographic B-Role
  • OGS Jumper Routing
  • COL D1 FD2 MSR
  • DOS3D PDP Check
  • IFM stowage Locate
  • Kubik 6 Belt Search
  • JPM1OF2 Lamp Housing Assembly (LHA) Replacement
  • National Engr Week Message
  • LHA JPM 1 of 2 R&R
  • P/TV Trash
  • PPS Wipe
  • Pre-gather COL A/L
  • Pre-gather Empty CTBs
  • SM PCS HD Swap Imaging
  • WINEARTH HISD Photos

Ground

  • Nominal Commanding
  • SSRMS Ops  [Complete]

Three-Day Look Ahead:

  • Saturday,12/21: USOS EVA 24
  • Sunday, 12,22:  EVA Tool Config, Procedure Review, Doug review
  • Monday, 12/23:  USOS EVA 25

QUICK ISS Status – Environmental Control Group:

                              Component Status
Elektron On
Vozdukh Manual
[СКВ] 1 – SM Air Conditioner System (“SKV1”) On
[СКВ] 2 – SM Air Conditioner System (“SKV2”) Off
Carbon Dioxide Removal Assembly (CDRA) Lab Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 Operate
Major Constituent Analyzer (MCA) Lab Off
Major Constituent Analyzer (MCA) Node 3 Operate
Oxygen Generation Assembly (OGA) Standby
Urine Processing Assembly (UPA) Standby
Trace Contaminant Control System (TCCS) Lab Off
Trace Contaminant Control System (TCCS) Node 3 Full up